OpenStudy (anonymous):

You are then asked to graph y = P(x) + 3. Explain how you would go about graphing this new function.

OpenStudy (anonymous):

so it's y-3 = P(x), in order for y-3 = P(x), y need to shift up 3 units

OpenStudy (anonymous):

Thank You